OpenStudy (anonymous):

rewrite the equation in vertex form name the vertex and y-intercept y=x^2-10x+15 A. y=(x-10)^2-10 vertext : (-10,-10) y-intercept: (0,10) B. y=(x-5)^2-10 vertex: (5,-10) y-intercept: (0,15) C. y=(x-5)^2 + 40 vertex: (5,-10) y-intercept: (0,15) D. y=(x-10)^2+20 vertex: (-10,-10) y-intercept:(0,-10)

OpenStudy (***[isuru]***):

y=x2−10+15 y=(x−5)2−25+10 y=(x−5)2−10 this is a graph with a minimum point to the function Y u will get the minimum value for the y when "x" is equal to 5 'cause when "x" is equal to 5 the term (x-5)^2 will be 0 .... and the function "y" will equal to -10 For any other value than 5 the term (x-5)^2 will take a positive value and therefor the value of function will increase.... So.. the "x" co-ordinate of the vertex is 5 and "y" cordinate of the vertex is -10 vertex = (5,-10) and u can find the "y" coordinate "y" intercept when the value of "x" is 0 which is 15 so... the intercept = (0,15) Hope this will help ya!!!!

OpenStudy (anonymous):

so it would be B

OpenStudy (***[isuru]***):

YEP !!!

OpenStudy (***[isuru]***):

ya got it right XD

OpenStudy (anonymous):

okay thanks lol :)

OpenStudy (***[isuru]***):

kk :-)